Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Bobbie Meng, Suzy Shaw, Leota WhitfieldAnnouncements
Ongoing Systems Updates
- No Lotto winners in Systems.
- Lots of equipment has gone out. PK Yonge picked up their stuff last week. Survey guys came by this morning & picked up quite a bit. Only 4 items were missed - their paperwork & ours did not jive. Will arrange for more Apple equipment to be picked up by PK Yonge.
- Replacement printer for Facilities has arrived. Scott has installed the cartridge for them. Printer needs to be checked in. The older, broken printer will need to be returned.
- CISCO routers will be donated to the SSRB lab instead of returning them to CISCO. When the paper work is finished on them, they'll be ready to go out.
- Bobbie will meet John Ashcraft & Daniel Cromwell at 1:30 PM today to supervise the pick up of 28 Cows bound for MSL.
- PC's to be ordered for System's software personnel will not have zip drives. Would only consider it if we had a need for backup that the server could not cover. The zip drives we have now are purely for public downloads.
- Need to set up another orientation session.
- We are unclear right now whether we will switch to Windows'98 before going to NT. Installing '98 over '95 is problematic, a mixture of both is a mess.
Software Projects Status
- Swapped out switch in West on Saturday. All switches now have web access software. Bill has created an experimental web page utilizing 'What's Up Gold' with activity display at 220.127.116.11. It is an online screen that opens a big chart. It also displays CORE routers.
- Microsoft is not yet finished with the patches for Outlook for client machines. The current danger to the mail system is over rated, the patches we've applied to the mail server seem to have stabilized the system.
- Winston has been through a crash course in Microsoft Access, helping Dan Roboussin to create a new database.
- ILL training session - IFM transaction reconciliation. Barbara Oliver didn't want Accounting to see the transactions ILL does through OCLC. Winston made the adjustments, ILL doesn't seem particularly happy with it. David Fuller is looking at other ILL software packages, the problem will be the interaction with the Accounting software.
- Personnel - permanent side is almost done.
- Special Collections is now entering data for their patron tracking.
- Orders dataset was not captured on Thursday morning. Dataset was in use, orders not created. Donna at FCLA was out, Maggie said it was OK to wait until today. FCLA recreated the dataset today, Sharon notified Charles McElroy. Talk ensued about a future migration to PC, Rich will check with FCLA.
- Primary key had to be reset on Evelyn's machine. NT4 had gone down & the pointer had to be reset. There's a patch to fix it, but when Winston tried to apply it, but it brought the whole system down.
- PAULA now has a drop down box for the acting ticket master.
- About 2 weeks ago, Dale came to Systems requesting information on patrons who had checked out ICPSR data in the past. With Sharon's help, Debra located the appropriate paperwork & forwarded it on to Dale.
- Debra completed the annual Library Statistics last week (again, with Sharon's assistance).
- Student Schedule for the Fall looks good, same group of students will be returning in two weeks.
- Discussion on voice mail messaging. We will try this in an attempt to get away from paper messages.
- Lee out this afternoon.
- Bill & Suzy probably out this Friday.
- Martha Hruska will be back on August 17.
- Lee out August 17-18.
- Bill out August 21 (Harris Corporation meeting).
- Debra out Friday, August 28.
- Lee out September 1.
- Labor Day September 7.
- Sharon out September 8.
- Lee out Sept. 17 - on vacation, then on to retirement.
Back to the Systems Homepage
Last updated February 24, 2000
by Debra Harris